A recent study shows that the annual cost of maintaining a building in Accra averages
200 Cedis with a variance of 260 Cedis. If a tax of 30% is introduced on all items
associated with the maintenance of building i.e. everything is made 30% more expensive.
Calculate the standard deviation of the annual cost of maintaining a building in Accra.
1
Expert's answer
2020-05-19T17:59:52-0400

"\text {Let Y represents the annual cost of maintaining}\\
\text { a building in Accra, then, E(Y)=200, V(Y)=260,}\\
\text {Let K represents the annual cost of maintaining}\\
\text { a building in Accra after adding a tax of 30\%,}\\
K=1.3Y\\
V(K)=V(1.3Y)=1.3^{2}V(Y)\\
=1.30^{2}(260)=439.4,\\
\therefore \sigma_{K}=\sqrt{V(K)}=\sqrt{439.4}=20.96"